OpenStudy (anonymous):

Solve by Factoring: x2 + 5x - 6 = 0

OpenStudy (solomonzelman):

x+y=5 x * y = -6 x and y are?

OpenStudy (studygurl14):

What two numbers add up to 5, and when multiplied, equals -6?

OpenStudy (studygurl14):

basically what @SolomonZelman put

OpenStudy (anonymous):

-3 and negative 2. I really don't know. i'm clueless. @StudyGurl14

OpenStudy (solomonzelman):

-3 + 2 = 5 ?!

OpenStudy (solomonzelman):

Since you are getting a positive middle term the negative number is less than the positive.

OpenStudy (studygurl14):

do this: What are the factors of -6? 2 and -3 -2 and 3 1 and -6 -1 and 6 right?

OpenStudy (studygurl14):

so what two factors from that list add up to 5?